# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v2

@@ III Design of Hamiltonians, ground states, and thermal states
In this section, we offer alternative interpretations and applications of the constrained energy minimization and constrained free energy minimization problems in (12) and (13), respectively. Namely, we show how applying the LMPW algorithms to these problems leads to the design of Hamiltonians with ground and thermal states satisfying fixed constraints on the expectation values of certain observables. In this way, the LMPW algorithms could be useful for the design of molecules and materials satisfying desired properties. Let us note that the design of ground states has been considered previously in the context of the density matrix renormalization group, in the special case that there is a single chemical potential that constrains a single conserved charge of the system [90, 91, 92].
@@ III.1 Design of ground states
The first equality in (71) follows by introducing $\mu$ as a vector of Lagrange multipliers to enforce the constraints in (69). The equality in (72) follows by moving the supremum to the outside, and the equality in (73) follows by simple algebra. The penultimate equality follows from the Sion minimax theorem for extended reals [93, Theorem 2.11]. The final equality follows by bringing in the minimum over all density matrices and recognizing that
@@
where $\lambda_{\min}(A)$ denotes the minimum eigenvalue of a Hermitian operator $A$. In the theory of semi-definite programming, the equality between the first line and the last line above is known as strong duality [63, Section 5.9.1].
The development in (71)–(75) demonstrates that, in the case that the constrained energy minimization problem in (70) is feasible (i.e., there exists at least one state satisfying the constraints in (69)), then its minimum value is finite and, for the dual optimization problem in (75), there exists an optimal choice of $\mu$, which we denote by $\mu^{*}$. Furthermore, for the optimal choice $\mu^{*}$, there is a corresponding Hamiltonian, $H-\mu^{*}\cdot Q$, such that there exists a state in its ground space satisfying the constraints in (69). To see this, let $\rho^{*}$ be a state that is optimal for (70), and let $\mu^{*}$ be optimal for (75). Then, the complementary slackness conditions of semi-definite programming imply that the following equality holds:
@@ V.1 Warm-starting the encoding of a single qubit
If $\left\|r\right\|=1$, then $\rho$ is a pure state, and if $\left\|r\right\|\leq 1$, then $\rho$ is a mixed state. Whenever $\rho$ is a mixed state, it can alternatively be written in terms of a chemical potential vector $\mu\coloneqq\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ and an inverse temperature parameter $\beta\geq 0$ as follows:
@@
The coordinates specified by $r$ are known as mixture coordinates, and those specified by $\mu$ and $\beta$ are known as exponential coordinates [95, Section 8.2]. The following theorem, proven in Appendix E, establishes a mapping from mixture coordinates to exponential coordinates for a single-qubit state:
@@ VI Simulation results
Figure 2: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the task of constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
In this section, we report the results of simulating the LMPW algorithms on several models of interest in quantum thermodynamics and the design of thermal states. Here, we focus on the one- and two-dimensional quantum Heisenberg models with a Hamiltonian having nearest- and next-to-nearest-neighbor interactions and the conserved, non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions (see Section II.3). We also report the results of simulations of a stabilizer thermodynamic system built from the perfect one-to-five qubit quantum error-correcting code.
@@
in order to track the performance of the algorithms. In (122), $E^{\star}$ is the true minimum energy output from a standard SDP solver and $\tilde{E}_{m}$ and $\tilde{\nabla}_{m}$ are the estimates of the energy and the gradient, respectively, at the $m$th step of the algorithm. If $\left|E^{\star}-\tilde{E}_{m}\right|$ is small, then we are guaranteed that the algorithm has output a good approximation of the true energy at the $m$th step, and if $\left\|\tilde{\nabla}_{m}\right\|$ is small, then, according to (18), we are guaranteed that the algorithm has found a solution that satisfies the constraints in (12) approximately. In all of our figures, we plot the logarithm of the error metric in (122), in order to make it easier to visualize the performance of the algorithms. In addition to using this error metric for plotting purposes, we also used the gradient norm $\left\|\tilde{\nabla}_{m}\right\|$ as the stopping criterion to determine algorithm convergence, terminating the optimization once it fell below a specified threshold.
@@
Our first set of simulations involves the two-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem from Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be $1$, 0, and $1$, respectively. Here we take the number of qubits to be fixed at four, arranged on a $2\times 2$ square lattice. Figure 2 compares the performance of all the LMPW algorithms, including the first- and second-order, classical and HQC algorithms, plotting the error metric in (122) versus the number of iterations. For the first- and second-order HQC algorithms, the average of five runs, along with one standard deviation shaded, is plotted. The figure demonstrates that all algorithms converge.
@@
Our next set of simulations reported here involve a stabilizer thermodynamic system constructed from the perfect one-to-five qubit code. Specifically, recall that a generating set for the code’s stabilizer consists of the following operators [56, Section 8.2]:
@@
Figure 4: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW classical algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
@@
Figure 5: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation. The task is constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions, and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW HQC algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Our final set of simulations involves the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem reviewed in Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Figure 4 demonstrates the performance of the first-order LMPW classical algorithm as the number of qubits in the model increases (specifically, the figure plots results for three, four, five, and six qubits). The figure plots the error metric in (122) versus the number of iterations, demonstrating that every case converges to the true value as the number of iterations increases.
@@ VI.1 Comparison of Standard Gradient Ascent and Nesterov’s Accelerated Method
Figure 6: The figure depicts the logarithm of the error metric in (122) versus the number of iterations for the LMPW classical first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Figure 7: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
@@
Figure 6 illustrates the performance of the standard and Nesterov-accelerated gradient ascent algorithms for the classical first-order optimization of the two-dimensional four-qubit Heisenberg model with nearest- and next-to-nearest-neighbor interactions. While both methods converge to the same solution, the Nesterov-accelerated method does so in significantly fewer iterations. Figure 7 presents the corresponding results for the HQC first-order optimization applied to the same task, again demonstrating that the Nesterov-accelerated method converges in significantly fewer iterations than the standard gradient ascent method.
@@ VII Conclusion
In this paper, we put forward several conceptual contributions, including framing constrained energy minimization as a method for designing ground states of Hamiltonians and constrained free energy minimization as a method for designing thermal states of Hamiltonians. We argued how the LMPW HQC algorithms can be used for designing these Hamiltonians and how this application has implications for designing molecules and materials with desired properties. We also introduced stabilizer thermodynamic systems as thermodynamic systems for which the Hamiltonian is constructed from the stabilizer operators of a stabilizer code and the conserved, non-commuting charges are constructed from the code’s logical operators. Then we demonstrated how the LMPW HQC algorithms can be understood as alternative methods for encoding quantum information into quantum error-correcting codes, which could be a useful approach whenever low-temperature thermal state preparation methods are experimentally accessible.
@@
The numerical contribution of our paper involved simulating the LMPW algorithms for several key thermodynamic systems of interest, including the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions and several stabilizer thermodynamic systems (the one-to-three qubit repetition code, the two-to-four qubit error-detecting code, and the perfect one-to-five qubit code). To test the performance of the LMPW algorithms on all of these examples, we employed standard SDP solvers to determine the true values of the constrained energy minimization problems. The LMPW classical algorithms can be understood as establishing alternative benchmarks for the HQC algorithms: the former algorithms act as noiseless simulators of the HQC algorithms, which focus exclusively on converging when measurement shot noise is not present, while our simulations of the latter incorporate measurement shot noise.
@@
Going forward from here, there are several avenues to pursue for future research. First, we think it would be worthwhile to employ matrix product states for simulating the thermal states needed in the LMPW classical algorithms, as used in prior work [96, 67, 97]. Very likely such an approach would lead an advanced classical algorithm and serve several purposes: first, it could be a useful algorithm on its own for understanding quantum thermodynamic systems; second, it could be helpful for solving or approximating the design of ground and thermal states of controllable Hamiltonians; third, it could be useful as a benchmark to set against the LMPW HQC algorithms. Regarding the latter point, any claimed quantum speedup for the LMPW HQC algorithms would have to outperform any possible classical algorithm for this task, and it is likely that approaches involving matrix product states would lead to the fastest possible classical algorithms.
@@ Appendix A Closeness of low-temperature thermal states and ground states of Hamiltonians
Now recall that $\Delta=\lambda_{2}-\lambda_{1}$. Then $e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\leq e^{\beta\Delta}$ for all $i\in\left\{2,\ldots,M\right\}$, which implies that

### v3

@@ III Design of Hamiltonians, ground states, and thermal states
In this section, we offer alternative interpretations and applications of the constrained energy minimization and constrained free energy minimization problems in (12) and (13), respectively. Namely, we show how applying the LMPW algorithms to these problems leads to the design of Hamiltonians with ground and thermal states satisfying fixed constraints on the expectation values of certain observables.
In this way, the LMPW algorithms could be useful for the design of molecules and materials satisfying desired properties. Specifically, they enable the precise targeting of specific electronic symmetry sectors (defined by particle number, spin, or other charges), which is a prerequisite for calculating the properties of molecular isomers and magnetic materials. As noted in recent literature regarding the Constrained Variational Quantum Eigensolver [22], standard unconstrained energy minimization is often insufficient for molecular design because the optimization can collapse to an incorrect sector (e.g., a neutral species instead of a cation). While penalty methods are commonly employed to enforce these constraints by energetically punishing unwanted sectors [23, 101, 24, 102], our framework offers a rigorous thermodynamic alternative that naturally targets specific sectors via chemical potentials, without requiring the manual tuning of penalty strengths. This approach connects directly to Reduced Density Matrix (RDM) theory [103], where constrained optimization is central to capturing electron correlation effects.
Let us note that the design of ground states has been considered previously in the context of the density matrix renormalization group, in the special case that there is a single chemical potential that constrains a single conserved charge of the system [104, 105, 106].
@@ III.1 Design of ground states
The first equality in (71) follows by introducing $\mu$ as a vector of Lagrange multipliers to enforce the constraints in (69). The equality in (72) follows by moving the supremum to the outside, and the equality in (73) follows by simple algebra. The penultimate equality follows from the Sion minimax theorem for extended reals [107, Theorem 2.11]. The final equality follows by bringing in the minimum over all density matrices and recognizing that
@@
where $\lambda_{\min}(A)$ denotes the minimum eigenvalue of a Hermitian operator $A$. In the theory of semidefinite programming, the equality between the first line and the last line above is known as strong duality [74, Section 5.9.1].
The development in (71)–(75) demonstrates that, in the case that the constrained energy minimization problem in (70) is feasible (i.e., there exists at least one state satisfying the constraints in (69)), then its minimum value is finite and, for the dual optimization problem in (75), there exists an optimal choice of $\mu$, which we denote by $\mu^{*}$. Furthermore, for the optimal choice $\mu^{*}$, there is a corresponding Hamiltonian, $H-\mu^{*}\cdot Q$, such that there exists a state in its ground space satisfying the constraints in (69). To see this, let $\rho^{*}$ be a state that is optimal for (70), and let $\mu^{*}$ be optimal for (75). Then, the complementary slackness conditions of semidefinite programming imply that the following equality holds:
@@ V.1 Warm-starting the encoding of a single qubit
If $\left\|r\right\|=1$, then $\rho$ is a pure state, and if $\left\|r\right\|<1$, then $\rho$ is a mixed state. Whenever $\rho$ is a mixed state, it can alternatively be written in terms of a chemical potential vector $\mu\coloneqq\left(\mu_{x},\mu_{y},\mu_{z}\right)\in\mathbb{R}^{3}$ and an inverse temperature parameter $\beta\geq 0$ as follows:
@@
The coordinates specified by $r$ are known as mixture coordinates, and those specified by $\mu$ and $\beta$ are known as exponential coordinates [109, Section 8.2]. The following theorem, proven in Appendix E, establishes a mapping from mixture coordinates to exponential coordinates for a single-qubit state:
@@ VI Simulation results
Figure 2: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the task of constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
In this section, we report the results of simulating the LMPW algorithms on several models of interest in quantum thermodynamics and the design of thermal states. Here, we focus on the one- and two-dimensional quantum Heisenberg models with a Hamiltonian having nearest- and next-nearest-neighbor interactions and the conserved, non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions (see Section II.3). We also report the results of simulations of a stabilizer thermodynamic system built from the perfect one-to-five qubit quantum error-correcting code.
@@
in order to track the performance of the algorithms. In (122), $E^{\star}$ is the true minimum energy output from a standard SDP solver (specifically, we employed CVXPY [110] for these benchmarks) and $\tilde{E}_{m}$ and $\tilde{\nabla}_{m}$ are the estimates of the energy and the gradient, respectively, at the $m$th step of the algorithm. If $\left|E^{\star}-\tilde{E}_{m}\right|$ is small, then we are guaranteed that the algorithm has output a good approximation of the true energy at the $m$th step, and if $\left\|\tilde{\nabla}_{m}\right\|$ is small, then, according to (18), we are guaranteed that the algorithm has found a solution that satisfies the constraints in (12) approximately. In all of our figures, we plot the logarithm of the error metric in (122), in order to make it easier to visualize the performance of the algorithms. In addition to using this error metric for plotting purposes, we also used the gradient norm $\left\|\tilde{\nabla}_{m}\right\|$ as the stopping criterion to determine algorithm convergence, terminating the optimization once it fell below a specified threshold.
@@
Our first set of simulations involves the two-dimensional quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem from Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be $1$, 0, and $1$, respectively. Here we take the number of qubits to be fixed at four, arranged on a $2\times 2$ square lattice. Figure 2 compares the performance of all the LMPW algorithms, including the first- and second-order, classical and HQC algorithms, plotting the error metric in (122) versus the number of iterations. For the first- and second-order HQC algorithms, the average of five runs, along with one standard deviation shaded, is plotted. The figure demonstrates that all algorithms converge.
@@
Our next set of simulations reported here involve a stabilizer thermodynamic system constructed from the perfect one-to-five qubit code. Specifically, recall that a generating set for the code’s stabilizer consists of the following operators [66, Section 8.2]:
@@
Figure 4: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW classical algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
@@
Figure 5: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation. The task is constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-nearest-neighbor interactions, and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW HQC algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Our final set of simulations involves the one-dimensional quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem reviewed in Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Figure 4 demonstrates the performance of the first-order LMPW classical algorithm as the number of qubits in the model increases (specifically, the figure plots results for three, four, five, and six qubits). The figure plots the error metric in (122) versus the number of iterations, demonstrating that every case converges to the true value as the number of iterations increases.
@@ VI.1 Comparison of Standard Gradient Ascent and Nesterov’s Accelerated Method
Figure 6: The figure depicts the logarithm of the error metric in (122) versus the number of iterations for the LMPW classical first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
Figure 7: The figure depicts the average of the logarithm of the error metric in (122) over five independent runs, plotted against the number of iterations, with shaded regions indicating one standard deviation, for the LMPW HQC first-order algorithm, comparing the standard gradient ascent method with Nesterov’s accelerated version. The task is constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively.
@@
Figure 6 illustrates the performance of the standard and Nesterov-accelerated gradient ascent algorithms for the classical first-order optimization of the two-dimensional four-qubit Heisenberg model with nearest- and next-nearest-neighbor interactions. While both methods converge to the same solution, the Nesterov-accelerated method does so in significantly fewer iterations. Figure 7 presents the corresponding results for the HQC first-order optimization applied to the same task, again demonstrating that the Nesterov-accelerated method converges in significantly fewer iterations than does the standard gradient ascent method.
@@ VII Conclusion
In this paper, we put forward several conceptual contributions, including framing constrained energy minimization as a method for designing ground states of Hamiltonians and constrained free energy minimization as a method for designing thermal states of Hamiltonians. We argued how the LMPW HQC algorithms can be used for designing these Hamiltonians and how this application has implications for designing molecules and materials with desired properties – specifically by enabling the precise targeting of specific electronic sectors (defined by particle number, spin, or other charges), which is a prerequisite for calculating the properties of molecular isomers and magnetic materials. We also introduced stabilizer thermodynamic systems as thermodynamic systems for which the Hamiltonian is constructed from the stabilizer operators of a stabilizer code and the conserved, non-commuting charges are constructed from the code’s logical operators. Then we demonstrated how the LMPW HQC algorithms can be understood as alternative methods for encoding quantum information into quantum error-correcting codes, which could be a useful approach whenever low-temperature thermal state preparation methods are experimentally accessible.
@@
The numerical contribution of our paper involved simulating the LMPW algorithms for several key thermodynamic systems of interest, including the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-nearest-neighbor interactions and several stabilizer thermodynamic systems (the one-to-three qubit repetition code, the two-to-four qubit error-detecting code, and the perfect one-to-five qubit code). To test the performance of the LMPW algorithms on all of these examples, we employed standard SDP solvers to determine the true values of the constrained energy minimization problems. The LMPW classical algorithms can be understood as establishing alternative benchmarks for the HQC algorithms: the former algorithms act as noiseless simulators of the HQC algorithms, which focus exclusively on converging when measurement shot noise is not present, while our simulations of the latter incorporate measurement shot noise.
Finally, we comment on the robustness of the LMPW algorithms in the low-temperature regime ($T\to 0$), which is required for high-precision ground state approximation. As discussed in Section III, the smoothness parameter of the dual objective function scales as $L\propto 1/T$. This implies that as the temperature decreases, the optimization landscape becomes stiffer, necessitating smaller step sizes ($\eta\propto T$) and leading to slower convergence rates. Additionally, for the second-order algorithms, the condition number of the Hessian increases, making the update steps more sensitive to sampling noise in the HQC setting. However, a key advantage of the LMPW framework is the guaranteed concavity of the dual landscape, which holds for all $T>0$. Unlike non-convex variational ansatzes that may get trapped in local minima as the system freezes, the LMPW algorithms are guaranteed to converge to the global optimum. Thus, while the computational cost (in terms of iterations and samples) increases at very low temperatures, the algorithms remain robust and do not suffer from the breakdown of convergence often observed in other variational methods.
Furthermore, our framework opens new avenues for exploring the interplay between thermodynamic geometry and quantum dynamics. For instance, the strictly concave geometry of the dual optimization landscape established in our work could provide a constructive method to probe phase transitions associated with the elimination of dynamical symmetries in systems with non-commuting charges [44]. Investigating how this geometry changes as non-commuting terms are introduced may offer deeper insights into why non-commuting charges can distinctly hinder or promote thermalization depending on the physical context.
@@
Going forward from here, there are several avenues to pursue for future research. First, we think it would be worthwhile to employ matrix product states for simulating the thermal states needed in the LMPW classical algorithms, as used in prior work [111, 78, 112]. Very likely such an approach would lead to an advanced classical algorithm and serve several purposes: first, it could be a useful algorithm on its own for understanding quantum thermodynamic systems; second, it could be helpful for solving or approximating the design of ground and thermal states of controllable Hamiltonians; third, it could be useful as a benchmark to set against the LMPW HQC algorithms. Regarding the latter point, any claimed quantum speedup for the LMPW HQC algorithms would have to outperform any possible classical algorithm for this task, and it is likely that approaches involving matrix product states would lead to the fastest possible classical algorithms.
@@ Appendix A Closeness of low-temperature thermal states and ground states of Hamiltonians
Now recall that $\Delta=\lambda_{2}-\lambda_{1}$. Then $e^{-\beta\left(\lambda_{i}-\lambda_{1}\right)}\leq e^{-\beta\Delta}$ for all $i\in\left\{2,\ldots,M\right\}$, which implies that